Week of 170227

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Kate (DB+CMS), Julia (WLCG), Maarten (ALICE), Marcelo (LHCb), Tomas (ATLAS), Nils (CM)
  • remote: Stephan (CMS), FaHui (ASGC), Xin (BNL), Matteo (CNAF), David (FNAL), Vincenzo (EGI), Victor (JINR), Sang-Un (KISTI), Preslav (KIT), Dmytro (NDGF), Andrew (NL-T1), Kyle (OSG), Pepe (PIC), John (RAL), Di (TRIUMF)

Experiments round table:

  • ATLAS reports ( raw view) -
    • BNL FTS server in troubles last week:
      • two alarm tickets released (closed):
      • after stable flow of transfers, alarm tickets were closed
      • remaining ggus ticket related to FTS activity shares:
    • pre-staging for reprocessing campaigns:
      • put large load on traffic within ATLAS grid
      • ~1.5 PB and 600k files still to be pre-staged.

  • CMS reports ( raw view) -
    • no major issues
    • Phase 1 and 2 upgrade (high pile-up) Monte Carlo generations in progress
    • To finalize the VO-feed switch we need a brief interrupt to the URL while the IP alias is being moved (during this week). Applications using curl to access the VO-feed need to add a "-L" to follow the redirect.

  • ALICE -
    • High activity
    • CNAF: CEs wrongly publishing zero waiting jobs since Fri morning (GGUS:126828)

  • LHCb reports ( raw view) -
    • Activity
      • MC Simulation, user analysis and data reconstruction jobs
    • Site Issues
      • T0:
        • Some settings were changed at EOS SRM and should have fixed last week's problem
        • Intervention on LHCb offline production database (LHCBR) to new hardware Wednesday 01/03/2017 from 10am to 12pm

Sites / Services round table:

  • ASGC:
  • BNL:
    • FTS and dCache went down last week, recovered shortly.
    • For on-going ATLAS staging from Tape, we tuned internal throttle in dCache, doubled the staging rate. More staging performance improvement is foreseen when upgrade to dCache 3.
  • CNAF:
  • EGI:
  • FNAL:
  • GridPP:
  • IN2P3:
  • JINR: All OK last week, but the farm is almost empty (jobs).
    Our real pledges as we have now (CMS): farm cores = 3200; SE Disk = 4000TB.
    SE MSS reconfigured and should not have problem for stage-in recent datasets.
    Our input data rate to Disk & MSS SE is relatively low mostly due to backlog at other side: T1_US_FNAL.*; T1_DE_KIT.*.
  • KISTI:
  • KIT: NTR
  • NDGF:
    • Three out of four CEs are down of different reasons (maintenance, filesystem problems, bad transfer system performance). Expecting one to be up today, and the second tomorrow.
  • NL-T1:
  • NRC-KI:
  • OSG: received ALARM tickets from ATLAS and redirected to BNL.
  • PIC: tape upgraded. Observed "double prestage" request. Will check with ATLAS if it is due to the prestage test.
  • RAL: There was a problem with the SAM SRM atlas tests at RAL over the weekend. This is not understood, but we believe production work was not affected.
  • TRIUMF: NTR

  • CERN computing services:
    • Batch farm rolling upgrade with kernel updates. Should be transparent.
    • Last week rebuild of 15% of batch capacity due to a hypervisor issue.
  • CERN storage services:
    • EOSCMS transparent intervention on the cluster headnode on March 7th from 7:30 to 8:30, clients will be stalled during the namespace failover although some interruption are to be expected.
  • CERN databases:
    • Network intervention on RAC52 switch tomorrow morning. Expected to be transparent. DBs at risk: Alice Stager, ADCR, ATONR
    • CMSONR ADG aliases will be moved to new HW tomorrow - no disconnection will be forced. Should be transparent.
    • CMSARC, LHCBR and PDBR database will be migrated to new HW tomorrow at 13:00/13:30. Up to 3 hours of downtime are expected.
    • LHCBR database will be migrated to new HW on Wednesday morning. Up to 2 hours of downtime are expected.
  • GGUS: NTR
  • Monitoring:
  • MW Officer: NTR
  • Networks:
  • Security: CVE-2017-6074:
    • EGI sent a heads-up, without any category ("expected" to be CRITICAL)
    • Trivial mitigation (disable dccp), should be blocked by SELinux

AOB:

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2017-02-27 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback